HUBBLE SPACE TELESCOPE DAILY REPORT #1862 PERIOD COVERED: 0000Z (UTC) 04/22/97 - 0000Z (UTC) 04/23/97 Daily Status Report as of 113/0000Z 1.0 ACTIVITIES SCHEDULED AND ACCOMPLISHED: 1.1 Completed Four Sets of FGS 7024 (FGS 3 Delta Mini-OFAD Post-SM2 = Calibration) FGS #3 was used to make astrometric observations of stars OF51, = OF63, OF89, OF91 and numerous other reference stars in order to reform a = mini-OFAD Post-SM2 calibration. The purpose of this activity is to = calibrate the changes to the optical field angle distortions of FGS unit = #3 that resulted from the servicing mission. The observations were = executed as scheduled, and no anomalies were noted. 1.2 Completed Five Sets of WF/PC-2 7553 (Pure Parallel Archive Survey = II) WF/PC-2 parallel observations were obtained in the ultraviolet in = search for interesting objects. This program is designed to find = extragalactic sources with near ultraviolet excesses (3000 A0) outside = the galactic plane and far ultraviolet sources (1600 A0) in the galactic = plane. The procedure was executed as scheduled, and no anomalies were = reported. 1.3 Completed NIC/3 7055 (NICMOS GRISM Validation) The Near Infrared Camera #3 was used to make observations of the = object G191-B2B in order to perform validation of the operations of the = three NICMOS grisms and supply the initial level of calibration for = spectroscopic observations. The observations were completed as planned = and no problems were reported. 1.4 Completed FGS 7025 (FGS/FGS Alignment Post-SM2 Calibration = (FGS1R)) The FGSs were used to make observations of the astrometric field = NGC-5617, and several other reference stars in order to work on the = FGS/FGS alignment during the post-SM-2 period calibration. These = observations were completed as scheduled, and no problems were reported. 1.5 Completed NIC/1/2 7043 (NICMOS Focus Monitor) The Near Infrared Cameras (Cameras 1, and 2) were used to make = observations of NGC-3603 in order to make focus calibration and = monitoring measurements. This proposal is built to run NIC/1/2 focus = sweeps on a weekly basis during SMOV. The observations were executed as scheduled, and no = anomalies were noted. 2.0 FLIGHT OPERATIONS SUMMARY: 2.1 Guide Star Acquisitions: Scheduled Acquisitions : 10 Successful : 10 Scheduled Reacquisitions : 7 Successful : 7 2.2 FHST Updates: Scheduled : 20 Successful : 20 2.3 Operations Notes: There were no significant operational problems reported. The battery #2 capacity test is still being executed. The STIS CCD detector was turned off at 112/2233Z. 3.0 SIGNIFICANT FORTHCOMING EVENTS: Continuation of normal science and SMOV operations. /DMH